# Planning without Search: Refining Frontier LLMs with Offline Goal-Conditioned RL

### v1

@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive performance across a wide range of reasoning and problem-solving tasks [38, 11, 43]. However, many complex tasks, such as goal-oriented dialogue, social interaction, and negotiation require LLMs to engage in complex multi-turn interactions, where the LLM agent might need to act strategically, anticipating the long-horizon outcomes of its decisions, while still adapting to unexpected responses. In such settings, an effective agent might take actions that are not immediately rewarding, but instead lay the ground for future success by either eliciting information or changing the interlocutor’s disposition. For example, an agent may ask clarifying questions to gather crucial information or attempt to persuade an interlocutor to shift their opinion, ultimately enabling the agent to achieve its objective. The space of possible outcomes when planning out such complex behaviors is enormous, and enabling such agents to act intelligently and strategically presents a major algorithmic challenge.
A traditional approach to equipping LLMs agents with such planning capabilities involves fine-tuning them using multi-turn reinforcement learning (RL) [33, 4]. While RL can in theory help LLMs optimize their behavior across extended interactions, the approach does not scale well to frontier models like GPT-4 [22] or Claude 3 [1]. In particular, RL algorithms are challenging due to their complexity to train and need for high number of samples, both of which make training large LLMs prohibitively expensive. Consequently, while RL has achieved some successes in multi-turn settings, its scalability limitations hinder broader adoption. Because of this, recent methods that leverage frontier models as LLM agents do not perform RL fine-tuning, but during inference, prompt the agent to simulate future trajectories to guide decision-making [44, 47].
@@
We further make a number of design decisions that provide for a computationally efficient and scalable method. First, rather than evaluating low-level utterances, we operate on high-level strategies, or thoughts, akin to chain-of-thought reasoning [38]. This abstraction significantly reduces the complexity of the decision space while preserving essential information. In this hierarchical design, our natural language critic assists the LLM agent in planning over high-level strategies, which can then be used to generate actions in the environment. By leveraging the critic, our approach also eliminates the need for any inference-time search, significantly improving its practicality on time-sensitive tasks. Instead of existing methods rely on feedback from inference-time search, we simply require evaluations by our natural language critic on both positive and negative hypothetical futures –or desirable and undesirable future occurrences to accomplishing the task at-hand, respectively.
@@ 2 Related Work
Large language models. Language models have shown impressive capabilities in text generation [6, 14, 8, 28, 40], translation [7], question answering [27], summarization [24, 39, 3], and code generation [5, 46]. Transforming LLMs into interactive agents capable of multi-step decision-making has been a recent focus of research. These embodied agents interact with environments by taking actions, observing responses, and adapting their strategies accordingly. Applications include robot learning  [32, 19], web navigation  [20], and text games involving dialogue  [41, 31, 47]. Our work falls in the realms of training better LLM agents by proposing a novel, more scalable approach.
@@
RL for LLMs. Reinforcement learning (RL) has been proposed to improve the performance of LLMs to much success. The most notable has been to use policy gradient methods to fine-tune LLMs from human or AI feedback  [2, 23, 36]. Such online methods have also been applied to train LLM agents in interactive settings, such as web search  [20] and embodied navigation  [45]. In contrast, offline RL methods have also been used to train LLM agents without the need for online samples  [34, 4]. However, RL methods require a high number of samples, and are notoriously difficult to train effectively; therefore, multi-turn RL has not been successfully scaled to frontier LLMs such as GPT-4 [22]. We also employ multi-turn RL, but we make several key innovations that make our approach much more scalable: (1) we consider RL over high-level strategies rather than environment actions; (2) we consider goal-conditioned RL where rather than learning values, we learn likelihoods of reaching particular goal states; and (3) we do not directly fine-tune an LLM, but rather learn a light-weight auxiliary value function that can be used to guide the search of the base LLM agent using only inference APIs. These innovations make our method practical to use for frontier LLMs.
@@
Reasoning and planning with LLMs. The latest frontier models have shown that strong reasoning and planning capabilities can be elicited by different prompting techniques  [38, 11]. Such techniques are central to the design of modern LLM agents. Notably, ReAct leverages chain-of-thought prompting [38] to allow LLM agents to generate a composite of reasoning and task-specific actions  [41]. However, ReAct relies on the base intuition of LLMs, which fails in more complex tasks such as web search or dialogue. More recent approaches have proposed self-supervised methods for LLM agents to enhance their reasoning capabilities  [31, 18, 47]. Reflexion adds a step of self-reflection to ReAct allowing LLM agents to refine their initial reasoning after some environment feedback  [31]. State-of-the-art methods employ tree-based search such as Monte Carlo tree search (MCTS) to guide the reasoning and planning of LLM agents [44, 47, 26]. These methods improve upon ReAct by leveraging the knowledge gained from search to improve the LLM agent’s reasoning capabilities. However, we posit that limited search is not sufficient for LLM agents to make data-driven decisions; existing approaches are limited by the amount of data agents can observe due to the prohibitive cost of inference-time search. In our work, we use offline RL to train value functions that can guide reasoning without the need for search during inference, resulting in efficient and data-driven agents. In contrast to explicit search, another popular paradigm for improving reasoning is self-refinement [9], where an LLM assesses its own generations intrinsically. Our method also involves refinement, but via an auxiliary goal-conditioned value function.
@@ 3 Preliminaries
Markov decision processes. We use the formalism of the Markov decision process (MDP), given by a tuple $M=({\mathcal{S}},{\mathcal{A}},P,r,\rho,\gamma)$, where ${\mathcal{S}}$ is the state space, ${\mathcal{A}}$ is the action space, $P$ is the transition function, $r$ is the reward function, $\rho$ is the initial state distribution, and $\gamma$ is the discount factor. When action $a\in{\mathcal{A}}$ is executed at state $s\in{\mathcal{S}}$, the next state is sampled $s^{\prime}\sim P(\cdot|s,a)$, and the agent receives reward $r$ with mean $r(s,a)$. Note that in many interactive settings, the state is actually partially observed, resulting in a POMDP instead of a MDP. For example, in dialogue, the state consists of the internal thoughts and intentions of other participants than the agent. However, prior analyses has shown that the two definitions can be made equivalent by instead defining states in the MDP as histories of observations so far, or in dialogue, the history of all utterances thus far [35].
LLM agents in MDPs. Tasks considered by LLM agents can be defined under this formalism as follows. At timestep $t$, the agent state $s_{t}$ of the MDP consists of the history of interaction thus far. In a dialogue, this will include the agent’s past utterances, as well as responses by all other participants. Following the ReAct agent by  Yao et al. [41], the agent action $a_{t}$ is actually composite, consisting of a thought $a_{t}^{\mathsf{tht}}$ where the agent performs a reasoning step, followed by the actual environment action $a_{t}^{\mathsf{env}}$. In dialogue, the thought could be the high-level strategy or plan the agent aims to execute, whereas the environment action is the actual utterance by the agent. Then, the transition function appends the environment action $a_{t}^{\mathsf{env}}$ as well as new observations by the environment $o_{t}$, such as responses by other interlocutors, to form the next state $s_{t+1}$.
@@
A policy can be extracted from the Q-function, i.e. $\pi^{*}(a|s)=\mathrm{1}[a=\arg\max_{a^{\prime}}Q^{*}(s,a^{\prime})]$. In offline RL, we want to estimate Q-function from a dataset $\mathcal{D}$ comprising of samples $(s,a,s^{\prime},r)$, where each sample corresponds to one timestep of interaction, and $a$ come from a behavior policy $\pi_{\beta}(\cdot|s)$ (which might correspond to a mixture of multiple policies). Kostrikov et al. [12] propose Implicit Q-learning (IQL), which trains Q-values $\widehat{Q}(s,a)$ and state-values $\widehat{V}(s)$ with the following loss functions:
$L_{Q}\!=\!\mathbb{E}_{(s,a,s^{\prime},r)\sim\mathcal{D}}\!\left[\left(r+\gamma\widehat{V}(s^{\prime})-Q(s,a)\right)^{2}\right]\,,\quad L_{V}\!=\!\mathbb{E}_{(s,a)\sim\mathcal{D}}\!\left[L_{2}^{\tau}\left(\widehat{Q}(s,a)-V(s,g)\right)\right]\,,$
@@ 4 PNLC: Planning with a Natural Language Critic
An overview of our approach can be found in Figure 2. During offline training, we aim to learn a goal-conditioned value function $Q(s,a^{\mathsf{tht}},g)$. Here, the goal $g$ is a possible future state that can result from following thought $a^{\mathsf{tht}}$ in state $s$, and $Q(s,a^{\mathsf{tht}},g)$ denotes the likelihood of reaching the goal future state $g$. Then, during inference, we can evaluate thoughts by an LLM agent by analyzing possible futures. Specifically, by sampling possible positive and negative future states and providing their likelihoods, an LLM agent can self-refine its original thought. We describe each stage in greater detail below, but defer specifics such as training hyperparameters and prompts used to Appendix A.
@@ 4.1 Training a Goal-Conditioned Value Function
Figure 2: During offline training, training samples are summarized and embedded, and a goal-conditioned value function (which is a light-weight 2 hidden-layer MLP) is trained over the embeddings. During inference, a natural language critic uses the value function to produce an informative analysis of possible future outcomes. This natural language value is used by the LLM agent to refine its proposed reasoning.
Our proposed natural language critic relies on a goal-conditioned value function to evaluate actions in lieu of inference-time search. As input, we require a dataset of task-specific trajectories $\mathcal{D}$ by any prior agent. The agent does not need to be near-optimal, but only exhibit a variety of different, even incorrect, strategies  [13]. In our experiments, we use an LLM agent with chain-of-thought prompting, which are known to exhibit suboptimal performance in tasks requiring long-term planning.
@@
The training samples are used to learn a goal-conditioned Q-value $\widehat{Q}(s,a^{\mathsf{tht}},g)$ by modifying the loss function of the IQL algorithm [12] to condition on goals:
@@
An illustration of the offline stage of our approach can be found in Figure 2. Note that though the summarization and embeddings require calls to an LLM, because they do not require particularly sophisticated reasoning, we are able to use cheaper LLM models to perform this step. In practice, we use GPT-3 [21] to extract embeddings, rather than GPT-4 [22] that is used in the final decision-making process. Also note that because our goal-conditioned Q-function does not take natural language inputs but rather embeddings, our learned Q-function is simply a MLP with 2 hidden layers rather than a full Transformer.
@@ 4.2 Planning using a Natural Language Critic
Now, we describe how our LLM agent leverages the goal-conditioned value function learned offline as a natural language critic for complex decision-making. Naively, the LLM agent could simply use the value function to take the action with the highest probability of reaching a desired goal. However, LLM agents historically perform better when allowed to refine their actions using feedback  [31, 18]. This is typically done post-hoc using trajectories from earlier attempts or from simulation. We instead use the natural language critic as a way to provide feedback over proposed actions, so that the LLM agent can refine its actions without the need to gather any trajectories during inference.
@@
The feedback provided by the natural language critic can be viewed as a compact summarization of a plethora of possible future trajectories, much more than can be obtained via inference-time search. Hence, the natural language critic can co-opt the traditional search procedure in traditional LLM reasoning processes. Namely, given a proposed thought and natural language value, the LLM agent is optionally allowed to suggest an improvement to its original thought, up to a total of $m$ iterations of refinement. In practice, we find that $m\!=\!2$ with only one additional round of refinement was sufficient to achieve good performance across our evaluated tasks. See Figure 2 for an illustration of the inference stage of our approach. We will show later that our method not only profits from efficiency but is also more effective than existing approaches across a variety of benchmarks.
@@ 5 Experiments
To demonstrate the effectiveness of PNLC, we evaluate our method on a variety of multi-turn LLM agent benchmark tasks: web shopping  [42], social deduction games  [16], and persuasion  [37].
@@ 5.1 Task Descriptions
Web Shopping. WebShop [41] is an online shopping environment where an agent processes unstructured text data (in the form of descriptions crawled from Amazon) to purchase a product given some initial user specifications. At the end, the agent receives a reward between 00 and $1$ depending on the similarity between the purchased and ground-truth desired item. The performance is gauged using two metrics: the score, reflecting the average reward received, and success rate, indicating the frequency with which the chosen product fulfills all specified conditions. All relevant methods train on the same dataset consisting of $12$k initial user instructions, of which we randomly held out $100$ for evaluation. With the remaining instructions, we generate a dataset of trajectories in which we simulate a suboptimal agent by prompting GPT-3.5 with few-shot examples similar to that done by  Yao et al. [41].
@@
Social Deduction. AvalonBench [16] is a testbed for LLM agents modeled after the team-based discussion game Resistance Avalon, where agents must jointly deceive other agents while deducing their roles, all through natural language dialogue. Each game consists of 5 players, 3 of which are on the good side and the remaining 2 evil. A notable asymmetry of this game is that all but one of the good players do not know other players’ identities, while evil players are aware of who their teammates are. The one exception is the role Merlin, which the LLM agent will play due to being the hardest role in the game. Merlin knows the identity of the evil players, and must guide the other good players whilst maintaining his own identity a secret; if his identity is discovered by the evil players, they are allowed to assassinate him, and the evil side wins the game. The LLM agent wins if the good side succeeds in at least 3 of 5 rounds, where a successful round means a subset of players chosen by a designated leader for the round all vote to accomplish a mission. Light et al. [16] implemented baseline agents for all roles in the game, which use GPT-3.5 [21] for discussion and a heuristic for action selection. We simulate 2.5k trajectories using these baseline agents, in which the baseline Merlin achieves a $21.4\%$ win rate.
@@
Persuasion. We consider a goal-oriented dialogue task, inspired by Wang et al. [37]. In this task, an LLM agent must persuade a participant to donate to Save the Children, a non-governmental organization dedicated to international assistance for children. At the end, the participant can choose to donate up to \$2 total. Because the donation amount is a very noisy signal, following prior work  [25], we model potential donors by prompting GPT-3.5 [21] to generate responses. To ensure that intelligent persuasion is required to complete the task successfully, the simulated donor is prompted to be skeptical of donating to charities. In our setup, the simulated user interacts with the LLM agent for up to 10 turns of dialogue, then must choose an amount to donate. We college 2.5k dialogues using naive prompting of GPT-3.5 [21] as the data collection agent. The average donation amount in the dataset is just \$$\$0.21$.

### v2

@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive performance across a wide range of reasoning and problem-solving tasks [wei2023chainofthought, kojima2022largelanguagemodelszeroshot, yao2023treethoughtsdeliberateproblem]. However, many complex tasks, such as goal-oriented dialogue, social interaction, and negotiation require LLMs to engage in complex multi-turn interactions, where the LLM agent might need to act strategically, anticipating the long-horizon outcomes of its decisions, while still adapting to unexpected responses. In such settings, an effective agent might take actions that are not immediately rewarding, but instead lay the ground for future success by either eliciting information or changing the interlocutor’s disposition. For example, an agent may ask clarifying questions to gather crucial information or attempt to persuade an interlocutor to shift their opinion, ultimately enabling the agent to achieve its objective. The space of possible outcomes when planning out such complex behaviors is enormous, and enabling such agents to act intelligently and strategically presents a major algorithmic challenge.
A traditional approach to equipping LLMs agents with such planning capabilities involves fine-tuning them using multi-turn reinforcement learning (RL) [snell2022offline, qtransformer]. While RL can in theory help LLMs optimize their behavior across extended interactions, the approach does not scale well to frontier models like GPT-4 [OpenAI2023GPT4] or Claude 3 [TheC3]. In particular, RL algorithms are challenging due to their complexity to train and need for high number of samples, both of which make training large LLMs prohibitively expensive. Consequently, while RL has achieved some successes in multi-turn settings, its scalability limitations hinder broader adoption. Because of this, recent methods that leverage frontier models as LLM agents do not perform RL fine-tuning, but during inference, prompt the agent to simulate future trajectories to guide decision-making [yu2023promptbasedmontecarlotreesearch, zhou2024languageagenttreesearch].
@@
We further make a number of design decisions that provide for a computationally efficient and scalable method. First, rather than evaluating low-level utterances, we operate on high-level strategies, or thoughts, akin to chain-of-thought reasoning [wei2023chainofthought]. This abstraction significantly reduces the complexity of the decision space while preserving essential information. In this hierarchical design, our natural language critic assists the LLM agent in planning over high-level strategies, which can then be used to generate actions in the environment. By leveraging the critic, our approach also eliminates the need for any inference-time search, significantly improving its practicality on time-sensitive tasks. Instead of existing methods rely on feedback from inference-time search, we simply require evaluations by our natural language critic on both positive and negative hypothetical futures –or desirable and undesirable future occurrences to accomplishing the task at-hand, respectively.
@@ 2 Related Work
Large language models. Language models have shown impressive capabilities in text generation [ghazvininejad-etal-2017-hafez, li2017learning, holtzman-etal-2018-learning, radford2019language, yang-klein-2021-fudge], translation [gu-etal-2017-trainable], question answering [pyatkin2022reinforced], summarization [paulus2017deep, wu2018learning, bohm-etal-2019-better], and code generation [codex, zhong2023study]. Transforming LLMs into interactive agents capable of multi-step decision-making has been a recent focus of research. These embodied agents interact with environments by taking actions, observing responses, and adapting their strategies accordingly. Applications include robot learning  [singh2022progpromptgeneratingsituatedrobot, min2022film], web navigation  [nakano2022webgpt], and text games involving dialogue  [yao2022react, shinn2023reflexionlanguageagentsverbal, zhou2024languageagenttreesearch]. Our work falls in the realms of training better LLM agents by proposing a novel, more scalable approach.
@@
RL for LLMs. Reinforcement learning (RL) has been proposed to improve the performance of LLMs to much success. The most notable has been to use policy gradient methods to fine-tune LLMs from human or AI feedback  [bai2022constitutional, ouyang2022training, touvron2023llama]. Such online methods have also been applied to train LLM agents in interactive settings, such as web search  [nakano2022webgpt] and embodied navigation  [zhai2024finetuninglargevisionlanguagemodels]. In contrast, offline RL methods have also been used to train LLM agents without the need for online samples  [snell2023offline, qtransformer]. However, RL methods require a high number of samples, and are notoriously difficult to train effectively; therefore, multi-turn RL has not been successfully scaled to frontier LLMs such as GPT-4 [OpenAI2023GPT4]. We also employ multi-turn RL, but we make several key innovations that make our approach much more scalable: (1) we consider RL over high-level strategies rather than environment actions; (2) we consider goal-conditioned RL where rather than learning values, we learn likelihoods of reaching particular goal states; and (3) we do not directly fine-tune an LLM, but rather learn a light-weight auxiliary value function that can be used to guide the search of the base LLM agent using only inference APIs. These innovations make our method practical to use for frontier LLMs.
@@
Reasoning and planning with LLMs. The latest frontier models have shown that strong reasoning and planning capabilities can be elicited by different prompting techniques  [wei2023chainofthought, kojima2022largelanguagemodelszeroshot]. Such techniques are central to the design of modern LLM agents. Notably, ReAct leverages chain-of-thought prompting [wei2023chainofthought] to allow LLM agents to generate a composite of reasoning and task-specific actions  [yao2022react]. However, ReAct relies on the base intuition of LLMs, which fails in more complex tasks such as web search or dialogue. More recent approaches have proposed self-supervised methods for LLM agents to enhance their reasoning capabilities  [shinn2023reflexionlanguageagentsverbal, madaan2023selfrefineiterativerefinementselffeedback, zhou2024languageagenttreesearch]. Reflexion adds a step of self-reflection to ReAct allowing LLM agents to refine their initial reasoning after some environment feedback  [shinn2023reflexionlanguageagentsverbal]. State-of-the-art methods employ tree-based search such as Monte Carlo tree search (MCTS) to guide the reasoning and planning of LLM agents [yu2023promptbasedmontecarlotreesearch, zhou2024languageagenttreesearch, putta2024agentqadvancedreasoning]. These methods improve upon ReAct by leveraging the knowledge gained from search to improve the LLM agent’s reasoning capabilities. However, we posit that limited search is not sufficient for LLM agents to make data-driven decisions; existing approaches are limited by the amount of data agents can observe due to the prohibitive cost of inference-time search. In our work, we use offline RL to train value functions that can guide reasoning without the need for search during inference, resulting in efficient and data-driven agents. In contrast to explicit search, another popular paradigm for improving reasoning is self-refinement [huang2023improve], where an LLM assesses its own generations intrinsically. Our method also involves refinement, but via an auxiliary goal-conditioned value function.
@@ 3 Preliminaries
Markov decision processes. We use the formalism of the Markov decision process (MDP), given by a tuple $M=({\mathcal{S}},{\mathcal{A}},P,r,\rho,\gamma)$, where $\mathcal{S}$} is the state space, $\mathcal{A}$} is the action space, $P$ is the transition function, $r$ is the reward function, $\rho$ is the initial state distribution, and $\gamma$ is the discount factor. When action $a\in{\mathcal{A}}$ is executed at state $s\in{\mathcal{S}}$, the next state is sampled $s^{\prime}\sim P(\cdot|s,a)$, and the agent receives reward $r$ with mean $r(s,a)$. Note that in many interactive settings, the state is actually partially observed, resulting in a POMDP instead of a MDP. For example, in dialogue, the state consists of the internal thoughts and intentions of other participants than the agent. However, prior analyses has shown that the two definitions can be made equivalent by instead defining states in the MDP as histories of observations so far, or in dialogue, the history of all utterances thus far [Timmer2010].
LLM agents in MDPs. Tasks considered by LLM agents can be defined under this formalism as follows. At timestep $t$, the agent state $s_{t}$ of the MDP consists of the history of interaction thus far. In a dialogue, this will include the agent’s past utterances, as well as responses by all other participants. Following the ReAct agent by  yao2022react, the agent action $a_{t}$ is actually composite, consisting of a thought $a_{t}^{\mathsf{tht}}$ where the agent performs a reasoning step, followed by the actual environment action $a_{t}^{\mathsf{env}}$. In dialogue, the thought could be the high-level strategy or plan the agent aims to execute, whereas the environment action is the actual utterance by the agent. Then, the transition function appends the environment action $a_{t}^{\mathsf{env}}$ as well as new observations by the environment $o_{t}$, such as responses by other interlocutors, to form the next state $s_{t+1}$.
@@
A policy can be extracted from the Q-function, i.e. $\pi^{*}(a|s)=\mathrm{1}[a=\arg\max_{a^{\prime}}Q^{*}(s,a^{\prime})]$. In offline RL, we want to estimate Q-function from a dataset $\mathcal{D}$ comprising of samples $(s,a,s^{\prime},r)$, where each sample corresponds to one timestep of interaction, and $a$ come from a behavior policy $\pi_{\beta}(\cdot|s)$ (which might correspond to a mixture of multiple policies). kostrikov2021offline propose Implicit Q-learning (IQL), which trains Q-values $\widehat{Q}(s,a)$ and state-values $\widehat{V}(s)$ with the following loss functions:
$L_{Q}\!=\!\mathbb{E}_{(s,a,s^{\prime},r)\sim\mathcal{D}}\!\left[\left(r+\gamma\widehat{V}(s^{\prime})-Q(s,a)\right)^{2}\right]\,,\qquad L_{V}\!=\!\mathbb{E}_{(s,a)\sim\mathcal{D}}\!\left[L_{2}^{\tau}\left(\widehat{Q}(s,a)-V(s,g)\right)\right]\,$,
@@ 4 PNLC: Planning with a Natural Language Critic
An overview of our approach can be found in Figure LABEL:fig:method. During offline training, we aim to learn a goal-conditioned value function $Q(s,a^{\mathsf{tht}},g)$. Here, the goal $g$ is a possible future state that can result from following thought $a^{\mathsf{tht}}$ in state $s$, and $Q(s,a^{\mathsf{tht}},g)$ denotes the likelihood of reaching the goal future state $g$. Then, during inference, we can evaluate thoughts by an LLM agent by analyzing possible futures. Specifically, by sampling possible positive and negative future states and providing their likelihoods, an LLM agent can self-refine its original thought. We describe each stage in greater detail below, but defer specifics such as training hyperparameters and prompts used to Appendix A.
@@ 4.1 Training a Goal-Conditioned Value Function
Figure 2: During offline training, training samples are summarized and embedded, and a goal-conditioned value function (which is just a small MLP) is trained over the embeddings.
Our proposed natural language critic relies on a goal-conditioned value function to evaluate actions in lieu of inference-time search. As input, we require a dataset of task-specific trajectories $\mathcal{D}$ by any prior agent. The agent does not need to be near-optimal, but only exhibit a variety of different, even incorrect, strategies  [kumar2022prefer]. In our experiments, we use an LLM agent with chain-of-thought prompting, which are known to exhibit suboptimal performance in tasks requiring long-term planning.
@@
The training samples are used to learn a goal-conditioned Q-value $\widehat{Q}(s,a^{\mathsf{tht}},g)$ by modifying the loss function of the IQL algorithm [kostrikov2021offline] to condition on goals:
@@
An illustration of the offline stage of our approach can be found in Figure 2. Note that though the summarization and embeddings require calls to an LLM, because they do not require particularly sophisticated reasoning, we are able to use cheaper LLM models to perform this step. In practice, we use GPT-3 [OpenAI2022ChatGPT] to extract embeddings, rather than GPT-4 [OpenAI2023GPT4] that is used in the final decision-making process. Also note that because our goal-conditioned Q-function does not take natural language inputs but rather embeddings, our learned Q-function is simply a MLP with 2 hidden layers rather than a full Transformer.
@@ 4.2 Planning using a Natural Language Critic
Figure 3: During inference, a natural language critic uses the value function to produce an informative analysis of possible future outcomes. This natural language value is used by the LLM agent to refine its proposed reasoning.
Now, we describe how our LLM agent leverages the goal-conditioned value function learned offline as a natural language critic for complex decision-making. Naively, the LLM agent could simply use the value function to take the action with the highest probability of reaching a desired goal. However, LLM agents historically perform better when allowed to refine their actions using feedback  [shinn2023reflexionlanguageagentsverbal, madaan2023selfrefineiterativerefinementselffeedback]. This is typically done post-hoc using trajectories from earlier attempts or from simulation. We instead use the natural language critic as a way to provide feedback over proposed actions, so that the LLM agent can refine its actions without the need to gather any trajectories during inference.
@@
The feedback provided by the natural language critic can be viewed as a compact summarization of a plethora of possible future trajectories, much more than can be obtained via inference-time search. Hence, the natural language critic can co-opt the traditional search procedure in traditional LLM reasoning processes. Namely, given a proposed thought and natural language value, the LLM agent is optionally allowed to suggest an improvement to its original thought, up to a total of $m$ iterations of refinement. In practice, we find that m=2m\!=\!2 with only one additional round of refinement was sufficient to achieve good performance across our evaluated tasks. See Figure 3 for an illustration of the inference stage of our approach. We will show later that our method not only profits from efficiency but is also more effective than existing approaches across a variety of benchmarks.
@@ 5 Experiments
To demonstrate the effectiveness of PNLC, we evaluate our method on a variety of multi-turn LLM agent benchmark tasks: web shopping  [yao2023webshop], social deduction games  [light2023avalonbenchevaluatingllmsplaying], and persuasion  [wang2019persuasion].
@@ 5.1 Task Descriptions
Web Shopping. WebShop [yao2022react] is an online shopping environment where an agent processes unstructured text data (in the form of descriptions crawled from Amazon) to purchase a product given some initial user specifications. At the end, the agent receives a reward between 0 and $1$ depending on the similarity between the purchased and ground-truth desired item. The performance is gauged using two metrics: the score, reflecting the average reward received, and success rate, indicating the frequency with which the chosen product fulfills all specified conditions. All relevant methods train on the same dataset consisting of 1212k initial user instructions, of which we randomly held out $100$ for evaluation. With the remaining instructions, we generate a dataset of trajectories in which we simulate a suboptimal agent by prompting GPT-3.5 with few-shot examples similar to that done by  yao2022react.
@@
Social Deduction. AvalonBench [light2023avalonbenchevaluatingllmsplaying] is a testbed for LLM agents modeled after the team-based discussion game Resistance Avalon, where agents must jointly deceive other agents while deducing their roles, all through natural language dialogue. Each game consists of 5 players, 3 of which are on the good side and the remaining 2 evil. A notable asymmetry of this game is that all but one of the good players do not know other players’ identities, while evil players are aware of who their teammates are. The one exception is the role Merlin, which the LLM agent will play due to being the hardest role in the game. Merlin knows the identity of the evil players, and must guide the other good players whilst maintaining his own identity a secret; if his identity is discovered by the evil players, they are allowed to assassinate him, and the evil side wins the game. The LLM agent wins if the good side succeeds in at least 3 of 5 rounds, where a successful round means a subset of players chosen by a designated leader for the round all vote to accomplish a mission. light2023avalonbenchevaluatingllmsplaying implemented baseline agents for all roles in the game, which use GPT-3.5 [OpenAI2022ChatGPT] for discussion and a heuristic for action selection. We simulate 2.5k trajectories using these baseline agents, in which the baseline Merlin achieves a $21.4\%$ win rate.
@@
Persuasion. We consider a goal-oriented dialogue task, inspired by wang2019persuasion. In this task, an LLM agent must persuade a participant to donate to Save the Children, a non-governmental organization dedicated to international assistance for children. At the end, the participant can choose to donate up to \$2 total. Because the donation amount is a very noisy signal, following prior work  [phuong2024evaluatingfrontiermodelsdangerous], we model potential donors by prompting GPT-3.5 [OpenAI2022ChatGPT] to generate responses. To ensure that intelligent persuasion is required to complete the task successfully, the simulated donor is prompted to be skeptical of donating to charities. In our setup, the simulated user interacts with the LLM agent for up to 10 turns of dialogue, then must choose an amount to donate. We college 2.5k dialogues using naive prompting of GPT-3.5 [OpenAI2022ChatGPT] as the data collection agent. The average donation amount in the dataset is just $\mathdollar 0.21$.